OpenStudy (anonymous):

Which line shows the first error in the solution? 8x-6=4x-2 (1) 8x-4x-6=4x-4x-2 (2) 4x-6=2 (3) 4x-6+6=2+6 (4) 4x=8 (5) 4x/4= 8/4 (6) x=2 (7) A. line 3 B. line 4 C. line 5 D. line 6

OpenStudy (anonymous):

@KirbyLegs

OpenStudy (anonymous):

Hey sorry about that.

OpenStudy (anonymous):

np

OpenStudy (anonymous):

i need help on this problem

OpenStudy (anonymous):

8x-6=4x-2 (1) 8x-4x-6=4x-4x-2 (2) 4x-6=2 (3) 4x-6+6=2+6 (4) 4x=8 (5) 4x/4= 8/4 (6) x=2 (7) Okay. Line one seems legit Line two is wrong. The 2 should be positive. Does that make sense?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so line two is wrong?>

OpenStudy (anonymous):

It doesn't look like you understand XD 8x-6=4x-2 Subtract 4x from each side (1) 8x-4x-6=4x-4x-2 (2) 4x-6=2 8x-4x-6 = 2 4x - 6 = 2 The sign of 2 does not change.

OpenStudy (anonymous):

lol im kinda confussed a little

OpenStudy (anonymous):

like how did yuo know it was wrong

OpenStudy (anonymous):

line two

OpenStudy (anonymous):

Simplify 8x-6 = 4x -2 Can you start with that?

OpenStudy (anonymous):

i think lol

OpenStudy (anonymous):

and im think no you cant start with that @KirbyLegs

OpenStudy (anonymous):

?

OpenStudy (anonymous):

the answer is line two/

OpenStudy (anonymous):

Do you understand why?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Okay that's good enough for me XD

OpenStudy (anonymous):

but line two is not an answer

OpenStudy (anonymous):

@KirbyLegs